|Operators ramped up the Main Injector last week.|
Operators and electrical engineering support personnel conducted Main Injector power supply testing and closed all of the Main Injector vacuum circuit breakers. Operators ramped the Main Injector up to 120 GeV. They also turned on all the NuMI power supplies and ramped them.
Linac personnel switched operations to Source B for the first time. It proved itself as efficient as Source A.
ANU, Main Injector and Recycler
Vacuum technicians completed the vacuum work and successfully leak-checked the beam pipes in Recycler section 40. They also completed the beam pipe installation at Recycler section 60.
- Service building low-conductivity water – IWS Inc. resealed the penetrations. However, the fabric bags need to be cut to length; the beads need to be repoured; and, after the beads are installed, they will need to be covered with foam to keep them in place.
- Power supply buss-in service buildings – Mechanical technicians refoamed the penetrations, but more work is yet needed.
- Empty RF buss – Riggers have installed the steel plates. The gravel will be poured next week.
- RF buss MI-60 – Technicians have assembled and installed all components.
NuMI personnel completed testing of Horn 1 thermocouples. Riggers dry-set R-blocks over Horn 1 and Target and moved the Target Hall shielding wall into place. Power supply personnel started pre-target power supply and magnet testing, which will continue until the next run begins. PPD personnel installed the controls rack for the Muon Monitor in Alcove 4 and the gas rack for Muon Alcove 4 and made the final connections for the Muon Alcove 4 gas system.
Alignment personnel aligned the B source and LEBT line. This is the section that gets moved into place when sources are switched. The three components are the source, trim magnet and solenoid. Alignment personnel also located the building corner points for the future MC-1 Building for FESS.
View the AD Operations Department schedule.